# Single vibronic level fluorescence spectra from Hagedorn wavepacket dynamics

### v2

@@ Abstract
In single vibronic level (SVL) fluorescence experiments, the electronically excited initial state is also excited in one or several vibrational modes. Whereas computing all contributing Franck-Condon factors individually becomes impractical in large systems, a time-dependent formalism has not been applied to simulate emission from arbitrary initial vibrational levels. Here, we use Hagedorn functions, which are products of a Gaussian and carefully generated polynomials, to represent SVL initial states. In systems where the potential is at most quadratic, the Hagedorn functions are exact solutions to the time-dependent Schrödinger equation and can be propagated with the same equations of motion as a simple Gaussian wavepacket. Having developed an efficient recursive algorithm to compute the overlaps between two Hagedorn wavepackets, we can now evaluate emission spectra from arbitrary vibronic levels using a single trajectory. We validate the method in two-dimensional global harmonic models by comparing it with quantum split-operator calculations. Additionally, we study the effects of displacement, distortion (squeezing), and Duschinsky rotation on SVL spectra. Finally, we demonstrate the applicability of the Hagedorn approach to high-dimensional systems on an example of displaced, distorted, and Duschinsky-rotated harmonic model with 100 degrees of freedom.
@@ I Introduction
Vibronic spectroscopy provides valuable insights into the quantum dynamics of molecules.[1] In single vibronic level (SVL) fluorescence spectrocopy, tuned narrow-band lasers are used to prepare a population of molecules excited to a specific vibronic level. The subsequent fluorescence decay of such a system can then be measured and provide important information about vibrational structures, excited-state relaxation, as well as the vibronic coupling between ground and excited states.[2, 3, 4, 5, 6, 7, 8]
@@
The simulation of SVL spectra of small molecules has been carried out using time-independent sum-over-state expressions,[9, 10, 11, 12, 13, 14, 15] but the computation of Franck-Condon factors in larger molecules becomes more challenging with the increasing number of vibrational states involved, especially due to the non-separable overlap integrals under Duschinsky mixing effects. A time-dependent approach is more efficient for larger systems and constitutes a more fundamental method that is not only naturally suited to simulate time-resolved and multidimensional spectroscopy[16, 17, 18, 19, 20] but can also readily incorporate Duschinsky rotation, Herzberg-Teller,[21, 22, 23, 24, 25, 26] and temperature effects.[27, 28, 29, 30] While time-dependent approaches have been applied to SVL[31, 32] and similar experiments,[33, 34, 35] their practical application has so far been confined to initial states that are excited at most to the first excited vibrational level and only in a single vibrational mode.
@@
For absorption or emission from the ground vibrational state, the thawed Gaussian approximation[36, 37, 38] (TGA) has been successfully applied to simulate vibrationally resolved electronic spectra in situations where a single Gaussian wavefunction is sufficient to adequately describe the dynamics.[39, 40, 41, 28, 18, 29, 42, 43] However, the SVL emission process involving a vibrationally excited initial state cannot be described by a single Gaussian wavepacket even at the initial time. Within the harmonic approximation, it may instead be described, at all times, by a Gaussian multiplied by a polynomial. Tapavicza’s efficient generating function approach[32] can treat initial states singly excited in one mode, whereas the extended TGA[44, 22, 28, 25, 45, 46] (ETGA) may accommodate simultaneous single excitations in multiple modes, but only in systems without Duschinsky rotation.[47] A more general and robust method is needed to address situations where the simultaneous excitation of multiple vibrational modes to higher levels is involved.
@@
By generalizing Dirac’s ladder operators for one-dimensional quantum harmonic oscillators to arbitrary dimensions, Hagedorn introduced a special pair of raising and lowering operators that generate a series of functions in the form of a Gaussian wavepacket multiplied by a polynomial.[48, 49, 50, 51, 52, 53] Unlike simple direct products of one-dimensional Hermite functions, Hagedorn functions are exact solutions to the time-dependent Schrödinger equation (TDSE) with an arbitrary many-dimensional harmonic potential when the parameters of the associated Gaussian are propagated with classical-like equations of motion. The Hagedorn functions also form a complete orthonormal basis that may be used to expand arbitrary numerically exact solutions to the TDSE with a general potential.[54] However, in this work, we consider exclusively Hagedorn wavepackets consisting of only one such Hagedorn function at all times. Henceforth, we will use “Hagedorn wavepackets” and “Hagedorn functions” synonymously.
@@
Here, we use the Hagedorn wavepacket dynamics to compute SVL emissions from different vibrational levels in two-dimensional harmonic models in order to validate the results against “exact” quantum split-operator calculations and, where applicable, (extended) TGA simulations. We also examine how nuclear displacement, mode distortion (squeezing), and Duschinsky rotation influence SVL spectra by analyzing the different spectra arising from three harmonic models that gradually incorporate these effects. One important advantage of our approach is that the propagation of Hagedorn wavepackets does not incur any additional computational cost in harmonic potentials beyond that of propagating a single Gaussian wavepacket. The Hagedorn approach is thus suitable for much higher dimensions than grid-based quantum methods. As a demonstration, we compute spectra in a two-state harmonic model system with 100 vibrational degrees of freedom.
@@ II Theory
Here, $\mu_{ge}$ is the electronic transition dipole moment, $\hat{H}_{g}$ the nuclear Hamiltonian associated with the ground electronic state, and $E_{e,K}=\hbar\omega_{e,K}$ the vibrational energy of the initial state. In this time-dependent framework, the initial wavepacket is evolved on the ground-state potential energy surface, and the overlaps between the initial wavepacket $|{K}\rangle$ and the propagated wavepacket $e^{-i\hat{H}_{g}t/\hbar}|{K}\rangle$ are computed in order to evaluate the spectrum.
@@
$\varphi_{0}[\Lambda_{t},S_{t}](q)=\frac{1}{(\pi\hbar)^{D/4}\sqrt{\det(Q_{t})}}\\ \times\exp\left(\frac{i}{\hbar}\left(\frac{1}{2}x^{T}\cdot P_{t}\cdot Q_{t}^{-1}\cdot x+p_{t}^{T}\cdot x+S_{t}\right)\right],$ (3)
with the shifted position $x:=q-q_{t}$, a phase factor $S_{t}$ related to the classical action, and a set of time-dependent parameters $\Lambda_{t}=(q_{t},p_{t},Q_{t},P_{t})$, where $q_{t}$ and $p_{t}$ represent the position and momentum of the wavepacket’s center, and $Q_{t}$ and $P_{t}$ are complex-valued $D$-dimensional matrices related to the position and momentum covariances.[69] The two matrices replace the “width matrix” $A_{t}=P_{t}\cdot Q_{t}^{-1}$ used in Heller’s parametrization[70, 53, 69] and satisfy the symplecticity conditions [52, 53]
@@
$Q_{t}^{\dagger}\cdot P_{t}-P_{t}^{\dagger}\cdot Q_{t}$ $=2i\mathrm{Id}_{D}.$ (5)
@@
where $\hat{x}:=\hat{q}-q_{t}$ and $\hat{y}:=\hat{p}-p_{t}$ are the shifted position and momentum operators. The ladder operators connect the thawed Gaussian wavepacket $\varphi_{0}$ to a family of functions in the form of a Gaussian multiplied by a polynomial, such that
@@
where $\langle j\rangle=(0,\dots,0,1,0,\dots,0)$ is the $j$-th unit vector in $D$ dimensions.[48, 51, 52, 53] In general, the polynomial factor cannot be expressed as a tensor product of univariate polynomials.[50, 71, 72, 53]
However, in the special case where $Q_{t}^{-1}\cdot\overline{Q_{t}}$ is diagonal, it can be represented as a tensor product of scaled Hermite polynomials.[53] If we adopt the mass-weighted normal-mode coordinates on the excited-state potential energy surface, the “position” matrix $Q_{0}$ of the initial state may be chosen to be diagonal with $Q_{0,jj}={\omega_{j}^{-1/2}}$, where $\omega_{j}$ is the angular frequency of the $j$-th vibrational mode. In this case, the initial Hagedorn function
$\varphi_{K}(q)=\langle q|K\rangle=\frac{\varphi_{0}(q)}{\sqrt{2^{|K|}K!}}\prod_{j=1}^{D}H_{K_{j}}\left(\sqrt{\frac{\omega_{j}}{\hbar}}\cdot x_{j}\right)$ (10)
@@
Notably, in a harmonic potential
@@
Hagedorn wavepackets, like thawed Gaussians, are exact solutions to the time-dependent Schrödinger equation. The parameters of the Gaussian associated with the Hagedorn wavepacket evolve according to the equations
@@
which are the same as the equations for propagating a Gaussian wavepacket.[51, 53, 69] Here, the subscript $t$ denotes the time, $m$ the real symmetric mass matrix (scalar in mass-weighted normal-mode coordinates), and $L_{t}$ is the Lagrangian. Remarkably, with Hagedorn wavepackets, the SVL emission from any vibrational level can be evaluated from the same trajectory $(\Lambda_{t},S_{t})$ at zero cost beyond propagating a simple Gaussian.
To finally obtain SVL spectra from the dynamics results, the evaluation of overlap integrals (2) between two Hagedorn wavepackets with different Gaussian parameters is needed. However, a straightforward analytical expression, like that available for overlaps between two Gaussian wavepackets, has remained elusive for Hagedorn wavepackets. Past applications of Hagedorn wavepackets require numerical quadrature for integrals, which becomes challenging in higher dimensions.[54, 60, 62, 63, 64, 65, 66] In Ref. 67, we have proposed instead a recursive algebraic scheme to compute these overlaps. This allows us to apply Hagedorn wavepackets in much higher dimensions, paving the way for their spectroscopic applications. In real molecular applications, the cost of computing the overlaps becomes negligible compared to the costly polyatomic electronic structure calculations, which only need to be carried out once for the central Gaussian trajectory, and do not have to be repeated for the vibrationally excited initial states.
@@
In this study, we focus solely on global harmonic models so that the results can be analyzed without additional approximation or errors. Whereas in many situations anharmonicity[24, 75, 76, 77] and non-adiabatic[78] effects must be taken into account, adiabatic harmonic potentials have served as a useful starting point for computing and assigning the spectra of many molecules.[79, 80, 9, 81, 82, 83, 34, 32, 84, 85, 35, 86]
@@ III.1 Two-dimensional harmonic potentials
We begin by considering two-dimensional model systems. They are the simplest multi-dimensional cases where, in the presence of the Duschinsky rotation effects, the dynamics of Hagedorn wavepackets differ from that of simple products of one-dimensional Hermite functions. Additionally, numerically exact quantum calculations are readily accessible for the verification of the surprisingly complicated spectra of these models.
We assume that the initial, excited-state electronic surface $V_{e}$ and the final, ground-state electronic surface $V_{g}$ can both be described by a quadratic function (11). In all cases, the ground-state surface $V_{g}$ is centered at $q_{\text{ref},g}=(0,0)$., and the initial wavepacket is centered at $q_{0}=q_{\text{ref},e}=(15,-15)$, i.e., at the minimum of the excited-state surface $V_{e}$, and has zero momentum $p_{0}=(0,0)$.
@@
Keeping $V_{e}$ the same in all systems, we then construct three different ground-state surfaces $V_{g}$ to demonstrate the effects of mode displacement, distortion, and Duschinsky rotation in a cumulative fashion. In each case, a Gaussian wavepacket, which gives rise to the emission spectrum from the ground vibrational level of $V_{e}$, is propagated for 40000 a.u. (20000 steps with a time step of 2 a.u.) on the surface $V_{g}$.
@@
For each system, we provide examples of emissions from various vibrational levels, including the ground vibrational level, a singly excited level, a level with higher excitations in one mode, and two levels with multi-mode excitations. Specifically, we simulate SVL spectra from initial vibrational levels $1^{a}2^{b}$, $(a,b)\in\{(0,0),(1,0),(0,3),(1,1),(2,1)\}$, where $a,b$ denote the vibrational quantum numbers in modes 1 and 2 of the excited electronic state.
@@
The autocorrelation function of the Hagedorn wavepacket associated with each initial vibrational level is computed every two steps using the overlap expressions derived in Ref. 67. The SVL emission spectra are then evaluated by Fourier transforming the autocorrelation functions multiplied with a Gaussian damping function, which results in a spectral broadening with a half-width at half-maximum of $50\,\text{cm}^{-1}$. Each spectrum is shifted so that the transition to the vibrational ground state ($1^{a}_{0}2^{b}_{0}$, where the subscripts denote the final vibrational quantum numbers in the ground electronic state after the transition) is at $0\,\text{cm}^{-1}$.
@@
Figure 1: Comparison of the $1^{2}2^{1}$ SVL spectra (upper panel) computed using quantum split-operator calculations (black line) and Hagedorn wavepacket dynamics (red dashes) in a displaced, distorted, and Duschinsky-rotated two-dimensional harmonic system, with the absolute difference between the two simulated spectra shown in blue in the lower panel.
@@
In order to validate the Hagedorn approach, we also propagated the initial states using the second-order Fourier split-operator algorithm with $256\times 256$ grid points. Figure 1 shows the comparison between the exact quantum and the Hagedorn dynamics results in the most complicated case treated here, i.e., the SVL spectrum involving a higher mixed initial excitation ($1^{2}2^{1}$) in the displaced, distorted, and Duschinsky-rotated system, which can be considered as the most comprehensive harmonic description of a real, anharmonic system. The spectrum simulated from Hagedorn dynamics is not only visually indistinguishable from the grid-based quantum calculation but the absolute differences are also extremely small (on the scale of $<10^{-11}$). The excellent agreement between the Hagedorn and exact quantum results is also observed for the other SVL spectra simulated in all three systems, which we now discuss in detail.
@@
Figure 2 contains the SVL emission spectra of the three systems computed using Hagedorn wavepackets. The spectral intensities in all spectra are scaled by the highest peak of the $1^{0}2^{0}$ spectrum in the displaced, distorted and rotated system because the other two systems can be thought of as simplified approximations of this most general harmonic system.
@@
Figure 2: Simulated SVL emission spectra of three two-dimensional two-level harmonic systems from five different initial vibrational levels (indicated by the labels on the right).
@@
If the ground- and excited-state surfaces were identical (up to a constant energy gap $v_{0}$), the resulting spectrum would consist of a single peak at a wavenumber corresponding to the electronic energy gap. Only transitions with $\Delta\upsilon=\mathbf{0}$ ($1^{a}_{a}2^{b}_{b}$) would be allowed since the vibrational wavefunctions with different vibrational quantum numbers $\upsilon$ in the two electronic states would be orthogonal to each other. The horizontal displacement between the two electronic surfaces breaks this symmetry and allows for more vibronic transitions.
@@
In the displacement-only system (first column in Fig. 2), the two displaced surfaces $V_{g}$ and $V_{e}$ have the same diagonal Hessian matrix $\kappa_{g}=\kappa_{e}$ corresponding to vibrational wavenumbers $\tilde{\nu}_{1}^{\prime\prime}=1100\,\text{cm}^{-1}$ and $\tilde{\nu}_{2}^{\prime\prime}=1800\,\text{cm}^{-1}$. Without Duschinsky rotation, the vibronic spectrum is the convolution of the spectrum in each vibrational degree of freedom. Consequently, the spectral peaks for transitions $1^{a}_{\alpha}2^{b}_{\beta}$ appear at $\alpha\tilde{\nu}_{1}^{\prime\prime}+\beta\tilde{\nu}_{2}^{\prime\prime}$, where $\alpha,\beta\in\mathbb{N}_{0}$.
@@
If the ground and excited electronic surfaces no longer have the same vibrational frequencies, the so-called “mode distortion” leads to changes in peak positions as well as in intensity (second column in Fig. 2) in addition to the effects from displacement. In the displaced and distorted system, the initial wavepackets are kept the same, but $V_{g}$ now has a different (although still diagonal) Hessian matrix $\kappa_{g,\text{distorted}}$ corresponding to vibrational wavenumbers $\tilde{\nu}_{1}^{\prime\prime}=750\,\text{cm}^{-1}$ and $\tilde{\nu}_{2}^{\prime\prime}=2200\,\text{cm}^{-1}$. As the ground electronic surface changes, the possible peak positions adapt accordingly to the combination of new frequencies, which can be observed by comparing the ordering and positions of the labeled peaks on the $1^{0}2^{0}$ spectra in the first and second columns. In this particular case, since $\tilde{\nu}_{2}^{\prime\prime}\approx 3\tilde{\nu}_{1}^{\prime\prime}$, the spectral peaks exhibit a more regular spacing, and certain broadened peaks, e.g. for $1^{0}_{3}2^{0}_{0}$ and $1^{0}_{0}2^{0}_{1}$ (labeled together on the $1^{0}2^{0}$ spectrum), actually overlap in this displaced and distorted system. The mode distortion also alters the intensity pattern, which is particularly evident in the $1^{1}2^{1}$ and $1^{2}2^{1}$ spectra (compare the first and second columns).

### v3

@@ Abstract
In single vibronic level (SVL) fluorescence experiments, the electronically excited initial state is also excited in one or several vibrational modes. Because computing such spectra by evaluating all contributing Franck-Condon factors becomes impractical (and unnecessary) in large systems, here we propose a time-dependent approach based on Hagedorn wavepacket dynamics. We use Hagedorn functions—products of a Gaussian and carefully generated polynomials—to represent SVL initial states because in systems whose potential is at most quadratic, Hagedorn functions are exact solutions to the time-dependent Schrödinger equation and can be propagated with the same equations of motion as a simple Gaussian wavepacket. Having developed an efficient recursive algorithm to compute the overlaps between two Hagedorn wavepackets, we can now evaluate emission spectra from arbitrary vibronic levels using a single trajectory. We validate the method in two-dimensional global harmonic models by comparing it with quantum split-operator calculations. Additionally, we study the effects of displacement, distortion (squeezing), and Duschinsky rotation on SVL fluorescence spectra. Finally, we demonstrate the applicability of the Hagedorn approach to high-dimensional systems on a displaced, distorted, and Duschinsky-rotated harmonic model with 100 degrees of freedom.
@@ I Introduction
Vibronic spectroscopy offers valuable insights into the quantum dynamics of molecules.[1] In single vibronic level (SVL) fluorescence spectroscopy, tuned narrow-band lasers are used to prepare a population of molecules excited to a specific vibronic level. Measuring the subsequent fluorescence decay of such a system can provide important information about vibrational structures, excited-state relaxation, as well as the vibronic coupling between ground and excited states.[2, 3, 4, 5, 6, 7, 8]
@@
The simulation of SVL spectra of small molecules has been carried out using time-independent sum-over-states expressions,[9, 10, 11, 12, 13, 14, 15] but the computation of Franck-Condon factors in larger molecules becomes more challenging with the increasing number of vibrational states involved, especially due to the non-separable overlap integrals under Duschinsky mixing effects. Methods based on the time-dependent picture are more efficient for larger systems and constitute a more fundamental approach that is naturally suited to simulating time-resolved and multidimensional spectroscopy.[16, 17, 18, 19, 20] Like the time-independent approach, the time-dependent one can also readily incorporate Duschinsky rotation, Herzberg-Teller,[21, 22, 23, 24, 25, 26] and temperature effects.[27, 28, 29, 30] Although the time-dependent approach has been applied to SVL[31, 32] and similar experiments,[33, 34, 35] its practical application has been confined to initial states excited at most to the first excited vibrational level and only in a single vibrational mode.
@@
In the case of absorption or emission from the ground vibrational state, the thawed Gaussian approximation[36, 37, 38] (TGA) has been successful in simulating vibrationally resolved electronic spectra in situations where a single Gaussian represents the evolving wavefunction reasonably well.[39, 40, 41, 28, 18, 29, 42, 43] However, the SVL emission process involving a vibrationally excited initial state cannot be described by a single Gaussian wavepacket even at the initial time. Within the harmonic approximation, it may instead be captured, at all times, by a Gaussian multiplied by a polynomial. Tapavicza’s efficient generating function approach[32] can treat initial states singly excited in one mode, whereas the extended TGA[44, 22, 28, 25, 45, 46] (ETGA) may accommodate simultaneous single excitations in multiple modes, but only in systems without Duschinsky rotation.[47] A more general method is needed for situations where the simultaneous excitation of multiple vibrational modes to higher levels is involved.
@@
By generalizing Dirac’s ladder operators for one-dimensional quantum harmonic oscillators to arbitrary dimensions, Hagedorn introduced a special pair of raising and lowering operators that generate a series of functions in the form of a Gaussian wavepacket multiplied by a polynomial.[48, 49, 50, 51, 52, 53] Unlike simple products of one-dimensional Hermite functions, Hagedorn functions are exact solutions to the time-dependent Schrödinger equation (TDSE) with an arbitrary many-dimensional harmonic potential when the parameters of the associated Gaussian are propagated with classical-like equations of motion. Hagedorn functions also form a complete orthonormal basis that may be used to expand arbitrary numerically exact solutions to the TDSE with a general potential.[54] However, in this work, we consider exclusively Hagedorn wavepackets consisting of only one Hagedorn function at all times. Henceforth, we use “Hagedorn wavepackets” and “Hagedorn functions” synonymously.
@@
Here, we first use the Hagedorn wavepacket dynamics to compute SVL emissions from different vibrational levels in two-dimensional harmonic models, where we can validate the results with “exact” quantum split-operator calculations. We also examine how nuclear displacement, mode distortion (squeezing), and Duschinsky rotation (mode mixing) influence SVL spectra by analyzing the different spectra arising from three harmonic models that gradually incorporate these effects. One important advantage of our approach is that the propagation of Hagedorn wavepackets does not incur any additional computational cost in harmonic potentials beyond that of propagating a single Gaussian wavepacket. The Hagedorn approach is thus suitable for much higher dimensions than grid-based quantum methods, and we demonstrate this by computing spectra of a two-state harmonic model system with 100 vibrational degrees of freedom.
@@ II Theory
Here, $\mu_{ge}$ is the electronic transition dipole moment, $\hat{H}_{g}$ is the nuclear Hamiltonian associated with the ground electronic state, $E_{e,K}=\hbar\omega_{e,K}$ is the total electronic and vibrational energy of the initial state, and the bar [e.g., in $\overline{C(t)}$] indicates the complex conjugate (we avoid using to prevent confusion with its use for the conjugate transpose in the mathematical literature on Hagedorn wavepackets). In this time-dependent framework, the initial wavepacket is evolved on the ground-state surface, and the overlaps between the initial wavepacket $|{K}\rangle$ and the propagated wavepacket $e^{-i\hat{H}_{g}t/\hbar}|{K}\rangle$ are computed in order to evaluate the spectrum.
@@
$\varphi_{0}[\Lambda_{t},S_{t}](q)=\frac{1}{(\pi\hbar)^{D/4}\sqrt{\det(Q_{t})}}\\ \times\exp\left[\frac{i}{\hbar}\left(\frac{1}{2}x^{T}\cdot P_{t}\cdot Q_{t}^{-1}\cdot x+p_{t}^{T}\cdot x+S_{t}\right)\right],$ (3)
with the shifted position $x:=q-q_{t}$, a phase term $S_{t}$ related to the classical action, and a set of time-dependent parameters $\Lambda_{t}=(q_{t},p_{t},Q_{t},P_{t})$, where $q_{t}$ and $p_{t}$ represent the position and momentum of the wavepacket’s center, and $Q_{t}$ and $P_{t}$ are complex-valued $D$-dimensional matrices related by[69]
$\operatorname{Cov}(\hat{q})=(\hbar/2)Q_{t}\cdot Q_{t}^{\dagger},\quad\operatorname{Cov}(\hat{p})=(\hbar/2)P_{t}\cdot P_{t}^{\dagger}$ (4)
to the position and momentum covariances. The two matrices replace the “width matrix” $A_{t}=P_{t}\cdot Q_{t}^{-1}$ used in Heller’s parametrization[70, 53, 69] and satisfy the symplecticity conditions [52, 53]
@@
$Q_{t}^{\dagger}\cdot P_{t}-P_{t}^{\dagger}\cdot Q_{t}$ $=2i\mathrm{Id},$ (6)
where $\mathrm{Id}$ is the $D$-dimensional identity matrix.
@@
where $\hat{x}:=\hat{q}-q_{t}$ and $\hat{y}:=\hat{p}-p_{t}$ are the shifted position and momentum operators. The ladder operators connect the thawed Gaussian wavepacket $\varphi_{0}$ to a family of Hagedorn functions $\varphi_{K}$ in the form of a Gaussian multiplied by a polynomial, such that
@@
where $\langle j\rangle=(\underbrace{0,\dots,0}_{j-1},1,\underbrace{0,\dots,0}_{D-j})$ is the $D$-dimensional unit vector with components $\langle j\rangle_{k}\equiv\delta_{jk}$.[48, 51, 52, 53] In Eqs. (9) and (10), the $j$-th components of $A^{\dagger}$ and $A$ respectively raise and lower the $j$-th component of the multi-index $K$ of the Hagedorn function $\varphi_{K}$ by a unity. In general, the polynomial factor cannot be expressed as a product of univariate polynomials.[50, 71, 72, 53]
However, in the special case where the inverse of $Q_{t}$ times the complex conjugate of $Q_{t}$ results in a diagonal $Q_{t}^{-1}\cdot\overline{Q_{t}}$ matrix, the polynomial factor of a Hagedorn function can be represented as a product of scaled Hermite polynomials $H_{n}$.[53] If we adopt the mass-weighted normal-mode coordinates of the excited-state potential energy surface, the “position” matrix $Q_{0}$ of the initial state may be chosen to be diagonal with $Q_{0,jj}={(m_{j}\omega_{j})^{-1/2}}$, where $\omega_{j}$ is the angular frequency of the $j$-th vibrational mode with a common mass $m_{j}=m$. In this case, the initial Hagedorn function
$\varphi_{K}(q)=\langle q|K\rangle=\frac{\varphi_{0}(q)}{\sqrt{2^{|K|}K!}}\prod_{j=1}^{D}H_{K_{j}}\left(\sqrt{\frac{m\omega_{j}}{\hbar}}\cdot x_{j}\right)$ (11)
@@
Notably, Hagedorn wavepackets, like thawed Gaussians, are exact solutions to the time-dependent Schrödinger equation in a harmonic potential
@@
which generally serves as a good approximation of the molecular potential energy surface near the equilibrium nuclear configuration. The parameters of the Gaussian associated with the Hagedorn wavepacket evolve according to the equations
@@
which are the same as the equations for propagating a purely Gaussian wavepacket.[51, 53, 69] Here, the subscript $t$ denotes the time, $m$ the real symmetric mass matrix (scalar in mass-weighted normal-mode coordinates), and $L_{t}$ is the Lagrangian.
To finally obtain SVL spectra from the dynamics results, the evaluation of overlap integrals (2) between two Hagedorn wavepackets with different Gaussian parameters is needed. However, a straightforward analytical expression, like that available for overlaps between two Gaussian wavepackets, has remained elusive for Hagedorn wavepackets. Past applications of Hagedorn wavepackets require numerical quadrature for integrals, which becomes challenging in higher dimensions.[54, 60, 62, 63, 64, 65, 66] In Ref. 67, we have instead proposed a recursive algebraic scheme to compute these overlaps. This allows us to apply Hagedorn wavepackets in much higher dimensions, paving the way for their spectroscopic applications.
Remarkably, Hagedorn wavepackets enable the evaluation of SVL emission spectra from any vibrational level using a single trajectory $(\Lambda_{t},S_{t})$ of the common guiding Gaussian. Therefore, regardless of the initial excitation, the electronic structure calculations (computationally much more expensive than the overlaps) need to be performed only once in real molecular applications.
@@
In this study, we focus solely on global harmonic models so that the results can be analyzed without additional approximation or errors. Whereas anharmonicity[24, 75, 76, 77] and non-adiabatic[78] effects must be taken into account in many situations, adiabatic harmonic potentials have served as a useful starting point for computing and assigning the spectra of many molecules.[79, 80, 9, 81, 82, 83, 34, 32, 84, 85, 35, 86]
@@ III.1 Two-dimensional harmonic potentials
We begin by considering two-dimensional model systems. They are the simplest multi-dimensional cases where, in the presence of Duschinsky rotation (mode mixing), the dynamics of Hagedorn wavepackets differ from that of simple products of one-dimensional Hermite functions. Additionally, numerically exact quantum calculations are readily accessible for the verification of the surprisingly complicated spectra of these models.
[…]
@@
Keeping $V_{e}$ the same in all systems, we construct three different ground-state surfaces $V_{g}$ to demonstrate the effects of mode displacement, distortion, and Duschinsky rotation in a cumulative fashion. In each case, a Gaussian wavepacket, which generates the emission spectrum from the ground vibrational level $|K=\mathbf{0}\rangle$ of $V_{e}$, is propagated for 40000 a.u. (20000 steps with a time step of 2 a.u.) on the surface $V_{g}$.
@@
For each system, we provide examples of emissions from the ground vibrational level, a singly excited level, a level with higher excitations in one mode, and two levels with multi-mode excitations. Specifically, we simulate SVL spectra from initial vibrational levels $1^{a}2^{b}$, $(a,b)\in\{(0,0),(1,0),(0,3),(1,1),(2,1)\}$, where $a,b$ denote the vibrational quantum numbers in modes 1 and 2 of the excited electronic state.
@@
The autocorrelation function of the Hagedorn wavepacket associated with each initial vibrational level is computed every two steps using the overlap expressions derived in Ref. 67. The SVL emission spectra are then evaluated by Fourier transforming the autocorrelation functions multiplied with a Gaussian damping function, which results in a spectral broadening with a half-width at half-maximum of $50\,\text{cm}^{-1}$. Each spectrum is shifted so that the transition to the vibrational ground level ($1^{a}_{0}2^{b}_{0}$, where the subscripts denote the final vibrational quantum numbers in the ground electronic state after the transition) is at $0\,\text{cm}^{-1}$.
@@
Figure 1: Comparison of the $1^{2}2^{1}$ SVL spectra computed using quantum split-operator calculations (black line) and Hagedorn wavepacket dynamics (red dots) in a displaced, distorted, and Duschinsky-rotated two-dimensional harmonic system, with the absolute difference between the two simulated spectra shown in blue in the bottom panel.
@@
To validate the Hagedorn approach, we also propagated the initial states using the second-order Fourier split-operator algorithm on a position grid ranging from $-128$ to $128$ in each dimension, with a total of $256\times 256$ equidistant grid points. Figure 1 shows the comparison between the exact quantum and the Hagedorn dynamics results in the most complicated case treated here, i.e., the SVL spectrum involving a higher mixed initial excitation ($1^{2}2^{1}$) in the displaced, distorted, and Duschinsky-rotated harmonic model system. This represents the most comprehensive description — within the harmonic oscillator model — of a molecular potential energy surface, which, in reality, has anharmonicity that the global harmonic dynamics considered in this work is not taking into account. The spectrum simulated from Hagedorn dynamics is not only visually indistinguishable from the grid-based quantum calculation but the absolute differences are also extremely small (on the scale of $<10^{-11}$). The excellent agreement between the Hagedorn and exact quantum results is also observed in the other simulated SVL spectra, which we now discuss in detail.
@@
Figure 2 contains the SVL emission spectra of the three systems computed using Hagedorn wavepackets. The spectral intensities in all spectra are scaled by the highest peak of the $1^{0}2^{0}$ spectrum in the displaced, distorted, and rotated system because the other two systems can be thought of as simplified approximations of this most general harmonic system.
@@
Figure 2: Simulated SVL emission spectra of three two-dimensional two-state harmonic systems from five different initial vibrational levels (indicated by the labels on the right).
@@
If the ground- and excited-state surfaces were not displaced, i.e., if they were identical (up to a constant energy gap $v_{0}$) with the same equilibrium geometry, the resulting spectrum would consist of a single peak at a wavenumber corresponding to the electronic energy gap. Only transitions with $\Delta v=\mathbf{0}$ ($1^{a}_{a}2^{b}_{b}$) would be allowed since the vibrational wavefunctions with different vibrational quantum numbers $v$ in the two electronic states would be orthogonal to each other. The horizontal displacement between the two electronic surfaces breaks this symmetry and allows for more vibronic transitions.
@@
In the displacement-only system (first column in Fig. 2), the two displaced surfaces $V_{g}$ and $V_{e}$ have the same diagonal Hessian matrix $\kappa_{g}=\kappa_{e}$ corresponding to vibrational wavenumbers $\tilde{\nu}_{1}^{\prime\prime}=1100\,\text{cm}^{-1}$ and $\tilde{\nu}_{2}^{\prime\prime}=1800\,\text{cm}^{-1}$. Without Duschinsky rotation, the vibronic spectrum is the convolution of the spectrum in each vibrational degree of freedom. Consequently, the peaks for transitions $1^{a}_{\alpha}2^{b}_{\beta}$ appear at $\alpha\tilde{\nu}_{1}^{\prime\prime}+\beta\tilde{\nu}_{2}^{\prime\prime}$, where $\alpha,\beta\in\mathbb{N}_{0}$.
@@
If the ground and excited electronic surfaces no longer have the same vibrational frequencies, the so-called “mode distortion” leads to changes in peak positions as well as in intensity (second column in Fig. 2) in addition to the effects from displacement. In the displaced and distorted system, the initial wavepackets are kept the same, but the force-constant matrix $\kappa_{g,\text{distorted}}$ defining the ground-state potential $V_{g}$ now yields vibrational wavenumbers $\tilde{\nu}_{1}^{\prime\prime}=750\,\text{cm}^{-1}$ and $\tilde{\nu}_{2}^{\prime\prime}=2200\,\text{cm}^{-1}$, different from those of the excited-state potential $V_{e}$. We still assume that the normal modes of the two electronic surfaces are aligned, and thus the matrix $\kappa_{g,\text{distorted}}$ remains diagonal. As the ground surface changes, the possible peak positions adapt accordingly to the combination of new frequencies, which can be observed by comparing the ordering and positions of the labeled peaks in the $1^{0}2^{0}$ spectra in the first and second columns. In this particular case, since $\tilde{\nu}_{2}^{\prime\prime}\approx 3\tilde{\nu}_{1}^{\prime\prime}$, the spectral peaks exhibit a more regular spacing, and certain broadened peaks, e.g. for $1^{0}_{3}2^{0}_{0}$ and $1^{0}_{0}2^{0}_{1}$ (labeled together on the $1^{0}2^{0}$ spectrum), actually overlap in this displaced and distorted system. The mode distortion also alters the intensity pattern, which is particularly evident in the $1^{1}2^{1}$ and $1^{2}2^{1}$ spectra (compare the first and second columns).